OpenStudy (lxelle):

Given that y = 1 when x = 0, solve the differential equation dy/dx = y(4-y), obtaining an expression for y in terms of x.

OpenStudy (perl):

dy/dx = y(4-y) dy /( y (4-y) ) = dx

OpenStudy (perl):

$$ \Large \int \frac{dy}{y (4-y)} = \int dx $$

OpenStudy (perl):

$$\Large \int \frac{dy}{y (4-y)} = \Large \int \frac{1}{4y} - \frac{1}{4(y-4)}~dy\\ $$

OpenStudy (lxelle):

Uh huh.

OpenStudy (lxelle):

Yup. Then?

OpenStudy (perl):

$$ \Large \int \frac{dy}{y (4-y)}= \int dx\\ \therefore \Large \int \frac{1}{4y} - \frac{1}{4(y-4)}~dy=\int dx \\ \therefore \Large \frac{1}{4}~ln|y| - \frac{1}{4}~ ln |y-4| = x+C\\ \therefore \Large \frac{1}{4}~ln|\frac{y}{y-4}| = x+C\\ \therefore \Large ln|\frac{y}{y-4}| = 4\cdot x+4C\\ \therefore \Large ln|\frac{y}{y-4}| = 4x+C^{*}\\[0.1 in] \therefore \LARGE e^{ \Large ln|\frac{y}{y-4}|} = e^{4x+C^{*}}\\ \therefore \Large\frac{y}{y-4} = e^{4x+C^{*}}\ $$

OpenStudy (perl):

C * is the same as a constant

OpenStudy (perl):

so we can just call it C

OpenStudy (lxelle):

Okay.

OpenStudy (perl):

$$ \Large \int \frac{dy}{y (4-y)}= \int dx\\ \therefore \Large \int \frac{1}{4y} - \frac{1}{4(y-4)}~dy=\int dx \\ \therefore \Large \frac{1}{4}~ln|y| - \frac{1}{4}~ ln |y-4| = x+C\\ \therefore \Large \frac{1}{4}~ln|\frac{y}{y-4}| = x+C\\ \therefore \Large ln|\frac{y}{y-4}| = 4\cdot x+4C\\ \therefore \Large ln|\frac{y}{y-4}| = 4x+C^{*}\\[0.1 in] \therefore \LARGE e^{ \Large ln|\frac{y}{y-4}|} = e^{4x+C^{*}}\\ \therefore \Large\frac{y}{y-4} = e^{4x+C}\\ \rm flip both sides \\ \therefore \Large\frac{y-4}{y} = \frac{1}{e^{4x+C}}\\ \therefore \Large \frac{y}{y} - \frac{4}{y} = \frac{1}{e^{4x+C}}\\ \therefore \Large 1 - \frac{4}{y} = \frac{1}{e^{4x+C}}\\ $$

OpenStudy (lxelle):

Okay.

OpenStudy (lxelle):

Do we put in the values alrd

OpenStudy (lxelle):

Nice.

OpenStudy (lxelle):

How do Im arrange?

OpenStudy (lxelle):

Thanks. I got it alrd. Could help men itch differentiation of y = x + cos2x?

OpenStudy (lxelle):

Is it1 - sin 2x?

OpenStudy (perl):

$$ \Large \int \frac{dy}{y (4-y)}= \int dx\\ \therefore \Large \int \frac{1}{4y} - \frac{1}{4(y-4)}~dy=\int dx \\ \therefore \Large \frac{1}{4}~ln|y| - \frac{1}{4}~ ln |y-4| = x+C\\ \therefore \Large \frac{1}{4}~ln|\frac{y}{y-4}| = x+C\\ \therefore \Large ln|\frac{y}{y-4}| = 4\cdot x+4C\\ \therefore \Large ln|\frac{y}{y-4}| = 4x+C^{*}\\[0.1 in] \therefore \LARGE e^{ \Large ln|\frac{y}{y-4}|} = e^{4x+C}\\ \therefore \Large|\frac{y}{y-4}| = e^{4x+C}\\ \therefore \Large |\frac{y}{y-4}| = e^{4x}\cdot e^C\\ \large \rm since~ e^{constant}= constant \\[0.1 in] \therefore \Large |\frac{y}{y-4}| = C ~e^{4x}\\ \text{ the sign of the absolute value} \pm \text{gets absorbed by the constant C } \\ \therefore \Large \frac{y}{y-4} = C ~e^{4x}\\ $$

OpenStudy (perl):

you need to use chain rule y = x + cos2x dy/dx = 1 - sin(2x) * 2